OpenStudy (anonymous):

which expression is equivelent to (³ √m)^7

OpenStudy (anonymous):

\[\left(3\sqrt{m}\right)^7=2187 m^{7/2} \]

OpenStudy (anonymous):

is that simplified cuz thats not an answer choice

OpenStudy (anonymous):

\[\left(\sqrt[3]{m}\right)^7=m^{7/3} \]

OpenStudy (anonymous):

wait nevermind it is

OpenStudy (anonymous):

but still dont understand how you got that

OpenStudy (anonymous):

\[\sqrt[3]{m}=m^{\frac{1}{3}} \]\[\left(m^{\frac{1}{3}}\right)^7=m^{7/3} \]